A 4.8 μC point charge is placed in an external uniform electric field that has a magnitude of 72000 N/C. At what distance from the charge is the net electric field zero?
1
Expert's answer
2020-09-17T14:29:29-0400

"E = \frac{k|q|}{r^2}"

The distance when the electric field is zero is given by

"E_1 = E_2"

"E_1 = \frac{kq}{r^2}"

"72000\; N/C= \frac{(9 \times 10^9)(4.8\times 10^{-6}C)}{r^2}"

"r^2 = 6"

"r = 2.45 \;m"

Answer: 2.45 m